### v3

@@ 3.1 Formal definition of the POf and theoretical properties
This section states the POf, given in Framework 1, and its properties enclosed in Theorem 1.
@@
Let $\mathbb{X}$ be a finite-dimensional normed space, and $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ be a partition such that the subproblem associated to Problem ($\mathbf{P}_{\texttt{i}}$) at $x$,
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{s}}(x)$)
admits a nonempty set $\Gamma(x)\subseteq\mathbb{Y}(x)$ of global solutions for all $x\in X\triangleq\{x\in\mathbb{X}:\mathbb{Y}(x)\cap Y\cap\Omega\neq\emptyset\}$, that we name the set of effective partition set indices. Define an oracle function $\gamma:X\to Y$, assumed tractable and satisfying $\gamma(x)\in\Gamma(x)$ for all $x\in X$. Define also the index function $\chi:\mathbb{Y}\to\mathbb{X}$ that, for all $y\in\mathbb{Y}$, maps $y$ to the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. Finally, define the problem of computing an optimal partition set index as the reformulation of Problem ($\mathbf{P}_{\texttt{i}}$),
@@
Let us stress that ensuring the tractability of the oracle function $\gamma$ is not within the scope of this work. In practice, we consider that we possess an optimization method to tackle Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) for any partition set index $x$. Then $\gamma$ consists in the function that, given any $x$, calls this method and returns its output. We state the POf in Framework 1 by assuming that the oracle function $\gamma$ computes, for all $x$, exactly a global solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$). However, we show in Sections 4.4 and 5.4 that the DFPOm remains efficient when $\gamma(x)$ computes an approximation of a global solution to Subproblem ($\mathbf{P}_{\texttt{s}}(x)$). The POf may even retain theoretical properties when $\gamma$ returns only an approximation of a local solution. We plan to formally study an alteration of Framework 1 allowing for this significant relaxation of the definition of $\gamma$ in a future work that we announce in Section 6.2. Similarly, how to find a relevant partition of $\mathbb{Y}$ in general is not within our scope either. Nevertheless, we believe that finding such a partition is often simple (although problem-dependent), as illustrated in Section 4 in a class of composite problems.
To study the properties of the POf, we define first a notion of generalized local solution. This notion, hinted in [3] but not formally stated, is suited for problems with a possibly nonclosed feasible set and a possibly not lower semicontinuous objective function. We introduce it in Definition 1 and discuss it in Remark 1, for a generic problem since we will apply it to either Problem ($\mathbf{P}_{\texttt{i}}$) and Reformulation ($\mathbf{P}_{\texttt{r}}$).
Definition 1 (Generalized local solutions).
@@
We may now claim the main property of the POf, Theorem 1: under Assumption 1, solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$) (either global, local, or generalized) yield solutions to Problem ($\mathbf{P}_{\texttt{i}}$) (of same nature). We discuss Assumption 1 in Remark 2, and we prove Theorem 1 in Section 3.3.
@@
    for all global solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{i}}$);
@@
    for all local solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$);
@@
    for all generalized local solutions $x^{*}\in\mathrm{cl}(X)$ to Reformulation ($\mathbf{P}_{\texttt{r}}$), the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
@@
Let us illustrate Definition 1 via two problems (with $\mathbb{A}\triangleq\mathbb{R}$) it is suited for. First, consider $f$ mapping any $a\in\mathbb{R}$ to $a+1$ if $a\geq 0$ and $\lvert{a}\rvert$ otherwise, and $A\triangleq\mathbb{R}$. Second, consider $f$ mapping any $a\in\mathbb{R}$ to itself, and $A\triangleq\mathbb{R}_{+}^{*}$. In both cases, the problem has no solution, but the point $a^{*}\triangleq 0$ has a remarkable property: the infimum of $\overline{f_{A}}$ is found as the limit of $(f(a^{k}))_{k\in\mathbb{N}}$ for some well-chosen sequence $(a^{k})_{k\in\mathbb{N}}$ converging to $a^{*}$. Definition 1 captures this property, and names $a^{*}$ a generalized solution (in these examples, it is a global one).
@@
Assumption 1 is likely nonstringent in practice. First, the index function $\chi$ is continuous when the partition set $\mathbb{Y}(x)$ varies continuously with its index $x\in\mathbb{X}$. This is controllable since the partition of $\mathbb{Y}$ is user-defined. Second, the partition of the set $X$ of effective partition set indices allows $\gamma$ to be piecewise uniformly continuous. This rejects only instances where $\gamma$ is locally unbounded, and is weaker than the Lipschitz continuity of $\gamma$ that is commonly assumed in parametric optimization.
3.2 The DFPOm and related convergence certifications
We now discuss our method to exploit the POf. It consists in solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using a DFO algorithm, so that we may apply Theorem 1 to the (exact) solution it returns. Accordingly, we name this method derivative-free partitioned optimization method (DFPOm). Reformulation ($\mathbf{P}_{\texttt{r}}$) is a usual blackbox optimization problem: its variables space $\mathbb{X}$ is likely low-dimensional (even when $\mathbb{Y}$ has infinite dimension), and, for all $x\in\mathbb{X}$, the objective value $\Phi(x)$ is computed on-the-fly by calling the oracle function $\gamma$ at $x$ and evaluating $\Phi(x)\triangleq\varphi(\gamma(x))$ (or $\Phi(x)\triangleq+\infty$ if $\gamma(x)$ is not defined). Then DFO algorithms are suited for Reformulation ($\mathbf{P}_{\texttt{r}}$), in particular those with a covering step [3] which ensures convergence to local solutions. Let us formalize the covering operator in Definition 2. For simplicity, we provide only its most concise version, but [3] also proposes more tractable variants.
@@ 3.2 The DFPOm and related convergence certifications
Given a point $x\in\mathbb{X}$ and a set $\mathcal{H}\subseteq\mathbb{X}$ acting as a history of past points evaluated, the covering operator seeks for a direction $d$ that yields a new point $x+d$ to evaluate as far as possible from $\mathcal{H}$ within a ball of constant radius (set to $1$ for ease of presentation) centred at $x$. The covering operator may be embedded into any DFO algorithm as an additional covering step. We formalize this framework as Algorithm 1. It consists in applying, at all iterations $k\in\mathbb{N}$, the covering operator at the current incumbent solutions $x^{k}$ and the current history $\mathcal{H}^{k}$ of past trial points. As $(x^{k})_{k\in\mathbb{N}}$ converges to some limit point $x^{*}$, this ensures that a dense set of points is evaluated near $x^{*}$, which allows asserting its local optimality. We state this claim in Proposition 1, a rewording of [3, Theorem 2] in our context.
@@
Assume that $\Phi$ is bounded below with bounded sublevel sets and that $X$ admits a partition into $N\in\overline{\mathbb{N}}^{*}$ ample continuity sets of $\Phi$. Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{r}}$). Moreover, if $x^{*}\in X$ and $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{r}}$).
We now state our second result. Theorem 2 formalizes the algorithmic DFPOm to solve Problem ($\mathbf{P}_{\texttt{i}}$), by solving Reformulation ($\mathbf{P}_{\texttt{r}}$) using any instance of Algorithm 1 and applying Theorem 1 to the returned solution. This result relies on the additional Assumption 2 that slightly refines Assumption 1, as discussed in Remark 3. The proof of Theorem 2 is given in Section 3.3.
@@
In Framework 1 under Assumptions 1 and 2, Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{r}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$, the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{i}}$). Moreover, if $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{i}}$).
As Theorem 2 shows, we only use Algorithm 1 to solve Reformulation ($\mathbf{P}_{\texttt{r}}$), a problem it is well suited for. Moreover, as Reformulation ($\mathbf{P}_{\texttt{r}}$) is a usual blackbox problem, most of the literature on DFO may be applied to it. Examples are discussed in Remark 4. Note also that the DFPOm is oblivious to whether the oracle function $\gamma$ returns exactly global solutions as requested in Framework 1. Theorem 2 describes only the case where Framework 1 is valid, but the DFPOm remains efficient in practice when $\gamma$ returns an approximation of a local solution, and a formal study of its convergence guarantees of the DFPOm in this case may be considered. We discuss these claims in Section 6.2, where we foresee a future work that will formally re-define the POf with a less stringent definition of $\gamma$ and study its properties. Finally, remark that Theorem 2 requires that Algorithm 1 runs ad infinitum. This is a usual drawback in DFO that, to our best knowledge, cannot be avoided. In practice, we may actually interrupt Algorithm 1 after some iteration index $\kappa\in\mathbb{N}$ (according to some stopping criterion), and consider the incumbent solution $x^{\kappa}$ as an approximation of an actual solution to Reformulation ($\mathbf{P}_{\texttt{r}}$). Then, in practice, the DFPOm returns $\gamma(x^{\kappa})$ as an approximation of a solution to Problem ($\mathbf{P}_{\texttt{i}}$).
We conclude this section with the practical instance of Algorithm 1 that we use in our numerical experiments in Sections 4 and 5. We instantiate $\mathcal{A}$ as the popular direct search method (DSM) [4, 6]. The resulting covering direct search method (cDSM) [3, Algorithm 1] is therefore a fit of the covering step into the DSM. We formalize the cDSM in Algorithm 2, and discuss its parameters in Remark 5.
Algorithm 2 cDSM [3, Algorithm 1] instance of Algorithm 1 considered in Sections 4 and 5.
@@
Assumption 2 builds additional mild requirements to Assumption 1. First, in practice, a partition of $\mathbb{Y}$ that makes the index function $\chi$ continuous usually yields its additional requirements as well. Second, the requirements on $\varphi_{|\gamma(X)}$ are standard, and moreover they are stated on $\varphi_{|\gamma(X)}$ instead of $\varphi$ directly (so even contexts where $\varphi$ has unbounded sublevel sets may be compatible). Finally, the additional requirements on the partition sets $(X_{i})_{i=1}^{N}$ inherited from Assumption 1 are a minimal safeguard against pathological behaviours [3] that are unlikely to appear in practice.
@@
The DFPOm is compatible with the notion of surrogate [9] from DFO. When $\gamma(x)$ is obtained, for all $x\in X$, by a numerical method solving ($\mathbf{P}_{\texttt{s}}(x)$), a surrogate is obtained by relaxing the precision of the method. It is also possible to order trial points according to a sensitivity analysis of Subproblem ($\mathbf{P}_{\texttt{s}}(x)$) with respect to $x$ via the generalized implicit function theorem [27]. This echoes concerns in parametric optimization about how the parameter influences the solution [18, 30].
@@
Sections 4 and 5 instantiate Algorithm 2 as follows. The point $x^{0}$ depends on the instance, while we set $\delta^{0}\triangleq 1$. We choose $(\lambda,\upsilon)\triangleq(\frac{1}{2},1)$ in Sections 4.1, 4.2, 4.3, 5.1 and 5.2, and $(\lambda,\upsilon)\triangleq(\frac{1}{2},2)$ in Section 5.3, and $(\lambda,\upsilon)\triangleq(\frac{3}{4},2)$ in Sections 4.4 and 5.4. The search step is skipped ($\mathcal{D}_{\tt{S}}^{k}\triangleq\emptyset$) and the poll step selects random orthogonal positive bases. The algorithm iterates until $\delta^{k}<1E^{-10}$.

### v5

@@ 3.1 Formal definition of the POf and theoretical properties
This section states the POf, given in Framework 1, and its properties, enclosed in Theorem 1.
@@
Let $\mathbb{X}$ be a finite-dimensional normed space, and $\mathbb{Y}=\sqcup_{x\in\mathbb{X}}\mathbb{Y}(x)$ be a partition such that the subproblem associated to Problem ($\mathbf{P}_{\texttt{ini}}$) at $x$,
$\underset{\begin{array}[]{c}y\in\mathbb{Y}(x)\end{array}}{\operatorname*{minimize}}\quad\varphi(y)\quad\mbox{subject to}\quad y\in\Omega,$ ($\mathbf{P}_{\texttt{sub}}(x)$)
admits a nonempty set $\Gamma(x)\subseteq\mathbb{Y}(x)$ of global solutions for all $x\in X\triangleq\{x\in\mathbb{X}:\mathbb{Y}(x)\cap Y\cap\Omega\neq\emptyset\}$, that we name the set of effective partition set indices. Define an oracle function $\gamma:X\to Y$, assumed tractable and satisfying $\gamma(x)\in\Gamma(x)$ for all $x\in X$. Define also the index function $\chi:\mathbb{Y}\to\mathbb{X}$ that, for all $y\in\mathbb{Y}$, maps $y$ to the index $x\in\mathbb{X}$ of the unique partition set $\mathbb{Y}(x)$ containing $y$. Finally, define the problem of computing an optimal partition set index as the reformulation of Problem ($\mathbf{P}_{\texttt{ini}}$),
@@
Let us stress that ensuring the tractability of the oracle function $\gamma$ is not within the scope of this work. In practice, we consider that we possess an optimization method to tackle Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) for any partition set index $x$. Then $\gamma$ consists in the function that, given any $x$, calls this method and returns its output. We state the POf in Framework 1 by assuming that the oracle function $\gamma$ computes, for all $x$, exactly a global solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). Our theoretical results are derived from this assumption. However, we show in Sections 5.4 and 6.4 that the DFPOm remains efficient when $\gamma(x)$ computes an approximation of a global solution to Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$). The POf may even retain theoretical properties when $\gamma$ returns only an approximation of a local solution. We announce in Section 7.2 a future work that will study an alteration of Framework 1 allowing for this relaxation of the definition of $\gamma$. Similarly, how to efficiently partition $\mathbb{Y}$ in general is not within our scope either. We believe that finding a relevant partition is often simple, as illustrated in an infinite-dimensional problem in Section 4 and a class of finite-dimensional composite problems in Section 5.
To study the POf, we define first a notion of generalized local solution. This notion, hinted in [3] but not formally stated, is suited for problems with a possibly nonclosed feasible set and a possibly not lower semicontinuous objective function. We introduce it in Definition 1 and discuss it in Remark 1, for a generic problem since we will apply it to either Problem ($\mathbf{P}_{\texttt{ini}}$) and Reformulation ($\mathbf{P}_{\texttt{ref}}$).
Definition 1 (Generalized local solution).
@@
We may now claim the main property of the POf, Theorem 1: under Assumption 1, solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$) (either global, local, or generalized) yield solutions to Problem ($\mathbf{P}_{\texttt{ini}}$) (of same nature). We discuss Assumption 1 in Remark 2, and we prove Theorem 1 in Section 3.3.
@@
    for all global solutions $x^{*}$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a global solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
@@
    for all local solutions $x^{*}\in X$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$);
@@
    for all generalized local solutions $x^{*}\in\mathrm{cl}(X)$ to Reformulation ($\mathbf{P}_{\texttt{ref}}$), the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
@@
Let us illustrate Definition 1 via two problems (with $\mathbb{A}\triangleq\mathbb{R}$) it is suited for. First, let $A\triangleq\mathbb{R}$ and $f:a\mapsto a+1$ if $a\geq 0$ and $\lvert{a}\rvert$ otherwise. Second, let $A\triangleq\mathbb{R}_{+}^{*}$ and $f:a\mapsto a$. In both cases, the problem has no solution, but the point $a^{*}\triangleq 0$ has a remarkable property: the infimum of $\overline{f_{A}}$ is found as the limit of $(f(a^{k}))_{k\in\mathbb{N}}$ for some well-chosen sequence $(a^{k})_{k\in\mathbb{N}}$ converging to $a^{*}$. Definition 1 captures this property, and names $a^{*}$ a generalized local solution (in these examples, it is a global one).
@@
Assumption 1 is likely nonstringent in practice. First, the index function $\chi$ is continuous when the partition set $\mathbb{Y}(x)$ varies continuously with its index $x\in\mathbb{X}$. This is controllable since the partition of $\mathbb{Y}$ is user-defined, and it holds when, e.g., the partition of $\mathbb{Y}$ is made by fixing some variables as in Sections 1.1 and 4. Second, the partition of $X$ allows $\gamma$ to be piecewise uniformly continuous. This rejects only instances where $\gamma$ is locally unbounded, and is weaker than the Lipschitz continuity of $\gamma$ that is commonly assumed in parametric optimization and bilevel optimization [16, 34].
3.2 The DFPOm and its related convergence certifications
We now discuss our method to exploit the POf. It consists in solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) using a DFO algorithm, and apply Theorem 1 to the (theoretically exact) solution it returns. Accordingly, we name this method derivative-free partitioned optimization method (DFPOm). Before going through this section, let us stress that we introduce the DFPOm for cases where solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) is challenging. The DFPOm is not needed when a solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$) is available via specific techniques, since Theorem 1 is compatible with any solution. Section 4 shows such an example, where Reformulation ($\mathbf{P}_{\texttt{ref}}$) is solved analytically. We also highlight that in practice, a DFO algorithm returns only an approximation of a solution. Yet, our numerical tests in Sections 5 and 6 prove that this discrepancy with our theoretical analysis does not alter the performance of the DFPOm.
Reformulation ($\mathbf{P}_{\texttt{ref}}$) is a usual blackbox optimization problem: its variables space $\mathbb{X}$ likely has a low dimension (even when $\mathbb{Y}$ has infinite dimension), and, for all $x\in\mathbb{X}$, the objective value $\Phi(x)$ is computed on-the-fly by calling the oracle function $\gamma$ at $x$ and evaluating $\Phi(x)\triangleq\varphi(\gamma(x))$ (or $\Phi(x)\triangleq+\infty$ if $\gamma(x)$ is not defined). Then DFO algorithms are suited for Reformulation ($\mathbf{P}_{\texttt{ref}}$), in particular those with a covering step [3] which ensures convergence to local solutions. Let us formalize the covering operator in Definition 2. We provide only one concise version, but [3] discusses more tractable variants.
@@ 3.2 The DFPOm and its related convergence certifications
Given a point $x\in\mathbb{X}$ and a set $\mathcal{H}\subseteq\mathbb{X}$ acting as a history of past points, the covering operator seeks for a direction $d$ that yields a new point $x+d$ as far as possible from $\mathcal{H}$ within a ball of unitary radius centred at $x$. The covering operator may be embedded into any DFO algorithm as an additional covering step that applies, at all iterations $k\in\mathbb{N}$, the covering operator at the current incumbent solution $x^{k}$ and the current history $\mathcal{H}^{k}$ of past trial points. We formalize this framework as Algorithm 1. As $(x^{k})_{k\in\mathbb{N}}$ converges to some point $x^{*}$, this ensures that a dense set of points is evaluated near $x^{*}$, which asserts its local optimality. We state this claim in Proposition 1, a rewording of [3, Theorem 2].
@@
Assume that $\Phi$ is bounded below with bounded sublevel sets and that $X$ admits a partition into $N\in\overline{\mathbb{N}}^{*}$ ample continuity sets of $\Phi$. Then Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$ are generalized local solutions to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Moreover, if $x^{*}\in X$ and $\Phi$ is lower semicontinuous at $x^{*}$, then $x^{*}$ is a local solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$).
We now state our second result. Theorem 2 formalizes the DFPOm to solve Problem ($\mathbf{P}_{\texttt{ini}}$), by solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) (exactly) using any instance of Algorithm 1 and applying Theorem 1 to the returned solution. This result relies on the additional Assumption 2 that slightly refines Assumption 1, as discussed in Remark 3. The proof of Theorem 2 is given in Section 3.3.
@@
In Framework 1 under Assumptions 1 and 2, Algorithm 1 solving Reformulation ($\mathbf{P}_{\texttt{ref}}$) generates a sequence $(x^{k})_{k\in\mathbb{N}}$ such that, for all $x^{*}\in\mathrm{acc}((x^{k})_{k\in\mathbb{N}})\neq\emptyset$, the set $\mathcal{ACC}(\gamma;x^{*})$ contains a generalized local solution to Problem ($\mathbf{P}_{\texttt{ini}}$). Moreover, if $x^{*}\in X$ and $\mathcal{ACC}(\gamma;x^{*})\subseteq\Omega$ and $\varphi$ is lower semicontinuous at all $y\in\mathcal{ACC}(\gamma;x^{*})$, then $\gamma(x^{*})$ is a local solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
As Theorem 2 shows, Algorithm 1 solves only Reformulation ($\mathbf{P}_{\texttt{ref}}$), a problem it is well suited for. Moreover, most of the literature on DFO is applicable to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Examples are discussed in Remark 4. Note also that the DFPOm is oblivious to whether the oracle function $\gamma$ returns exactly global solutions as requested in Framework 1. Theorem 2 describes only the case where Framework 1 is valid, but Sections 5 and 6 show that the DFPOm remains efficient when $\gamma$ returns an approximation of a local solution. We also foresee in Section 7.2 a future work that will formally re-define the POf with a less stringent definition of $\gamma$. Finally, remark that Theorem 2 requires for Algorithm 1 to run ad infinitum. This is a usual drawback in DFO that, to our best knowledge, cannot be avoided. In practice, we interrupt Algorithm 1 according to some stopping criterion, at some iteration index $\kappa\in\mathbb{N}$, and we return the incumbent solution $x^{\kappa}$ as an approximation of a solution to Reformulation ($\mathbf{P}_{\texttt{ref}}$). Then the DFPOm returns $\gamma(x^{\kappa})$ as an approximation of a solution to Problem ($\mathbf{P}_{\texttt{ini}}$).
We conclude this section with the practical instance of Algorithm 1 that we use in our numerical experiments in Sections 5 and 6. We instantiate $\mathcal{A}$ as the popular direct search method (DSM) [4, 6]. The resulting covering direct search method (cDSM) [3, Algorithm 1] is therefore a fit of the covering step into the DSM. We formalize the cDSM in Algorithm 2, and discuss its parameters in Remark 5.
Algorithm 2 cDSM [3, Algorithm 1] instance of Algorithm 1 considered in Sections 5 and 6.
@@
Assumption 2 builds additional mild requirements to Assumption 1. First, in practice, a partition of $\mathbb{Y}$ that makes the index function $\chi$ continuous usually yields its additional requirements as well. For example, it is the case in all problems from the class we study in Sections 5 and 6. Second, the requirements on $\varphi_{|\gamma(X)}$ are standard, and moreover they are stated on $\varphi_{|\gamma(X)}$ instead of $\varphi$ directly (so even contexts where $\varphi$ has unbounded sublevel sets may be compatible). Finally, the additional requirements on the partition sets $(X_{i})_{i=1}^{N}$ inherited from Assumption 1 are a minimal safeguard against pathological behaviours [3] that are unlikely to appear in practice.
@@
The DFPOm is compatible with the notion of surrogate [9] from DFO. When $\gamma(x)$ is obtained, for all $x\in X$, by a numerical method solving ($\mathbf{P}_{\texttt{sub}}(x)$), a surrogate is obtained by relaxing the precision of the method. It is also possible to order trial points according to a sensitivity analysis of Subproblem ($\mathbf{P}_{\texttt{sub}}(x)$) with respect to $x$ via the generalized implicit function theorem [30]. This echoes concerns in parametric optimization about how the parameter influences the solution [19, 34].
@@
Sections 5 and 6 instantiate Algorithm 2 as follows. The point $x^{0}$ depends on the instance, while we set $\delta^{0}\triangleq 1$. We choose $(\lambda,\upsilon)\triangleq(\frac{1}{2},1)$ in Sections 5.1, 5.2, 5.3, 6.1 and 6.2, and $(\lambda,\upsilon)\triangleq(\frac{1}{2},2)$ in Section 6.3, and $(\lambda,\upsilon)\triangleq(\frac{3}{4},2)$ in Sections 5.4 and 6.4. The search step is skipped ($\mathcal{D}_{\tt{S}}^{k}\triangleq\emptyset$) and the poll step selects random orthogonal positive bases. The algorithm iterates until $\delta^{k}<1E^{-10}$.